# MFC 5.0: An exascale many-physics flow solver

### v3

@@ 4.1.5 Euler–Lagrange sub-grid bubble dynamics
The Euler–Lagrange model for sub-grid bubble dynamics is based on the volume-averaged equations of motion and describes the dynamics of a mixture of dispersed bubbles in a compressible liquid. Mixture properties $\overline{(\cdot)}$ are defined as ()¯$\overline{(\cdot)}=(1-\alpha)(\cdot)_{l}+\alpha(\cdot)_{g}$, where $\alpha$ is the volume fraction of the gas or void fraction, and the subscripts $l$ and $g$ denote the liquid and gas phase, respectively. Assuming zero slip-velocity between the liquid and gas phase and applying the volume averaging yield the following set of inhomogeneous equations
@@
The volumetric oscillations of the sub-grid bubbles due to pressure variations in the surrounding medium are modeled using the Keller–Miksis equation of section 4.1.4. The Lagrangian field equation is integrated using a fourth-order accurate Runge–Kutta scheme with adaptive time-stepping.
@@
where $N_{\mathrm{bub}}$ is the number of bubbles, and $V_{n}$ is the volume of bubble $n$. We use the continuous, second-order, truncated Gaussian function for the kernel. With $\alpha$, all the terms on the right-hand side of 13 can be computed. More details on the numerical implementation can be found in Maeda and Colonius [9] and Fuster and Colonius [48].
@@ 4.1.6 Fluid–elastic structure interaction
Fluid–structure interaction with hypoelastic materials is modeled by adding the elastic shear stress $\bm{T}^{e}$ to the Cauchy stress tensor of the five-equation and six-equation models described in section 2.1.1 and section 2.1.2, respectively. The model relies on transforming strains to strain rates using the Lie objective temporal derivative [49, 50]. The strain rates are computed using high-order accurate central finite differences. The mixture energy is also updated to include the elastic contribution, which gives
@@
where $\bm{l}$ is the velocity gradient, and for an isotropic Kelvin–Voigt material $\dot{\bm{T}}^{e}=2G\bm{D}^{d}$ [51] are used to derive a time evolution equation for the elastic stresses. Here, the deviatoric component of the velocity gradient is $\bm{D}^{d}=\bm{D}-\mathrm{tr}(\bm{D})\bm{I}/3$. Appending this evolution equation to the five-equation model gives the five-equation model with hypoelasticity
@@
$\frac{\partial\rho\bm{u}}{\partial t}+\nabla\cdot(\bm{u}\otimes\bm{u}+p\bm{I}-\bm{T}^{e})$ $=\nabla\cdot\bm{T}^{v},$
@@
Details of the hypoelastic model implementation can be found in Spratt [52].
Fluid–structure interaction with hyperelastic materials is modeled with an evolution equation for the reference map using the reference map technique (RMT) of [53]. The gradient of the reference map $\bm{\xi}$, the inverse of the deformation gradient $\bm{F}$,
@@
and the elastic deviatoric contribution to the Cauchy stress tensor $\bm{T}^{e}$ are computed. For a compressible neo-Hookean material model, the elastic deviatoric part of the Cauchy stress tensor is
@@
where $J$ is the determinant of $\bm{b}$ and $\bm{I}$ is the identity tensor and the hyperelastic energy is $e^{e}=G\left(I_{b}-3\right)/2$, where $I_{b}$ is the first invariant of $\bm{b}$. Further details of the hyperelastic model implementation can be found in Barbosa et al. [54].
@@ 4.1.7 Chemical reactions and combustion
High-fidelity simulations of reacting flows are critical to designing efficient propulsion systems. Such simulations require models for the effects that chemical reactions have on the flow. There are two major differences with respect to simulations of inert flow. First, mixture composition varies locally due to chemical reactions, and as a result, fluid properties, like viscosity, do as well. This effect must be accounted for to achieve acceptable accuracy; it is typically achieved using transport equations for individual species. Second, because chemical reactions release heat, temperature changes can become substantial, so variable thermodynamic properties are needed for accuracy. As we expand in what follows, we now incorporate these effects through detailed models in the gas phase.
@@
Next, we provide detailed expressions for the chemical source term. Their implementation uses code generation and is discussed in section 4.3.7. We simulate gas-phase combustion by integrating section 4.1.7 along with the equations for conservation of momentum and total energy density section 2.1.1. Plans to extend this capability to multi-phase combustion with liquid or solid fuels entail the implementation of detailed expressions for phase changes [55, 56, 57].
@@
where $A_{n},\,b_{n},\,T_{a,n}$ are the pre-exponential factor, temperature exponent, and activation temperature of the $n^{\mathrm{th}}$ reaction. The equilibrium constant in 16 is determined from standard equilibrium thermodynamics; detailed expressions can be found elsewhere [58, 59].
As stated, the effects of combustion on mixture thermodynamics must also be accounted for. For the equation of state, we assume a mixture of ideal gases, so
@@
where $\{\hat{h}_{m}(T)\}_{m=1}^{N_{\mathrm{species}}}$ are the species molar enthalpies, modeled using NASA polynomials [60]:
@@
with fit coefficients $\{\hat{C}\}_{r=0}^{5}$ determined from experiments and collision integrals [61]. The fitting procedure for 22 includes heats of formation so, in concert with 20 and 21, it accounts for combustion heat release. In practice, to obtain the temperature, 20 is solved iteratively using a standard Newton method.
[…]
Figure 3: (a) Temperature and (b) velocity profiles of a one-dimensional reactive shock tube compared to the results of Martínez-Ferrer et al. [62].
@@ 4.1.8 Surface tension at diffuse material interfaces
Surface tension for two-fluid flows is implemented using the model of Schmidmayer et al. [64]. An advection equation for the color function $c$, which is 00 in one fluid and $1$ in the other, is added to the system of equations. This color function is then used to calculate the capillary stress tensor
@@
$\frac{\partial\rho\bm{u}}{\partial t}+\nabla\cdot(\rho\bm{u}\bm{u}+p\bm{I}+\bm{\Omega}-\bm{T}^{v})$ $=0,$
@@
where $\varepsilon_{0}$ is the capillary mixture energy. The six-equation model with surface tension is conservative, except for the pressure relaxation terms, and follows the second law of thermodynamics. This model has been validated against analytic solutions to the Laplace pressure jump and experimental results for shock droplet interaction [64].
@@ 4.1.9 Magnetohydrodynamics
Magnetohydrodynamics (MHD) governs the behavior of conducting fluids under the influence of magnetic fields. The governing equations for ideal MHD are [65]
@@
Here, $p$ is the thermal pressure and $\mathbf{B}$ is the magnetic field. The total energy density is $E=\rho e+\frac{1}{2}\rho|\mathbf{u}|^{2}+\frac{1}{2}|\mathbf{B}|^{2}$.
@@ 4.1.10 Relativistic Magnetohydrodynamics
$\mathbf{U}=\begin{pmatrix}D\\[2.5pt] \mathbf{m}\\[2.5pt] \tau\\[2.5pt] \mathbf{B}\end{pmatrix}:=\begin{pmatrix}\Gamma\rho\\[2.5pt] \Gamma^{2}\rho h\mathbf{u}+|\mathbf{B}|^{2}\mathbf{u}-(\mathbf{u}\cdot\mathbf{B})\mathbf{B}\\[2.5pt] \Gamma^{2}\rho h-p+\frac{1}{2}\left[|\mathbf{B}|^{2}+|\mathbf{u}|^{2}|\mathbf{B}|^{2}-(\mathbf{B}\cdot\mathbf{u})^{2}\right]-\Gamma\rho\\[2.5pt] \mathbf{B}\end{pmatrix}.$
Here, $\rho$ denotes the rest-mass density, $\Gamma$ is the Lorentz factor, $p$ is the thermal pressure, and $h=1+e+p/\rho$ is the specific enthalpy. The rest-mass contribution is subtracted from the total energy density to reduce numerical errors at low velocities.
The governing equations are [66]:
@@
$\frac{\partial\mathbf{m}}{\partial t}+\nabla\cdot\left[\mathbf{u}\otimes\mathbf{m}-\frac{1}{\Gamma^{2}}\mathbf{B}\otimes\mathbf{B}-(\mathbf{u}\cdot\mathbf{B})\,\mathbf{u}\otimes\mathbf{B}+p\mathbf{I}\right]$ $=0,$
@@
Recovery of the primitive variables from these highly nonlinear conservative variables requires an iterative scheme. In this work, we solve a single nonlinear equation using the Newton-Raphson method [66].
@@ 4.2.1 General characteristic boundary conditions
The priority for time-dependent boundary conditions for hyperbolic equations requires a smooth egress of the outgoing artifacts with a minimal effect on the interior solution. For this purpose, the non-reflecting characteristic boundary conditions [32] are used, which performs a characteristic decomposition of the governing equation at the boundary. Ignoring the contribution of the viscous and transverse terms, this is
@@
where $q_{c}$ is the set of conservative variables, $\Lambda_{x}=\operatorname{diag}(\lambda_{j})=\operatorname{diag}(u-c,u,u,u,u+c)$ are the eigenvalues of the Jacobian, and $R_{x}$ and $L_{x}$ are the right and left eigenvectors. The contribution of the characteristics can be further decomposed into incoming and outgoing waves depending on the sign of the eigenvalue as
@@
with $v$ and $w$ being the transverse velocities. The generalized characteristic boundary conditions make use of the solution outside the domain at ghost points ($q_{p}^{\mathrm{(g)}}$) to incorporate the contribution of the incoming characteristics [67]. The characteristics $\bm{\mathcal{L}}$ in this case are
@@
where $\Delta$ is the grid spacing between the boundary ($q_{p}$) and the ghost point ($q_{p}^{\mathrm{(g)}}$). In practice, $\Delta$ is set to be the grid spacing between the boundary and the first interior point for accuracy and stability. For a subsonic outflow at the right boundary, coefficients $\mathcal{L}^{i}_{2}$ through $\mathcal{L}^{i}_{5}$ are set to 00 as these correspond to outgoing waves. The incoming characteristic $\mathcal{L}^{i}_{1}$ is calculated as

### v4

@@ 4.1.5 Euler–Lagrange sub-grid bubble dynamics
The Euler–Lagrange model for sub-grid bubble dynamics is based on the volume-averaged equations of motion and describes the dynamics of a mixture of dispersed bubbles in a compressible liquid. Mixture properties $\overline{(\cdot)}$ are defined as $\overline{(\cdot)}=(1-\alpha)(\cdot)_{l}+\alpha(\cdot)_{g}$, where $\alpha$ is the volume fraction of the gas or void fraction, and the subscripts $l$ and $g$ denote the liquid and gas phase, respectively. Assuming zero slip-velocity between the liquid and gas phase and applying the volume averaging, yields the following set of inhomogeneous equations
@@
The volumetric oscillations of the sub-grid bubbles due to pressure variations in the surrounding medium are modeled using the Keller–Miksis equation of section 4.1.4. The Lagrangian field equation is integrated using a fourth-order accurate Runge–Kutta scheme with adaptive time stepping.
@@
where $N_{\mathrm{bub}}$ is the number of bubbles, and $V_{n}$ is the volume of bubble $n$. We use the continuous, second-order, truncated Gaussian function for the kernel. With $\alpha$, all the terms on the right-hand side of 12 can be computed. More details on the numerical implementation can be found in maeda2018eulerian and fuster2011modelling.
[…]
Figure 5: Evolution of an isolated bubble in response to a single cycle of a sinusoidal pressure wave. The analytical solution of the Keller–Miksis equation was reported by maeda2018eulerian.
[…]
Figure 6: Pressure profiles at the origin of the bubble screen using the Euler–Lagrange and Euler–Euler subgrid models. Mono and polydisperse (log-normally distributed with $\sigma_{p}=0.3$) bubble clouds are tested as labeled.
@@ 4.1.6 Fluid–elastic structure interaction
Fluid–structure interaction with hypoelastic materials is modeled by adding the elastic shear stress $\bm{T}^{e}$ to the Cauchy stress tensor of the five-equation and six-equation models described in section 2.1.1 and section 2.1.2, respectively. The model relies on transforming strains to strain rates using the Lie objective temporal derivative [Guilaume2015, guillaume2016]. The strain rates are computed using high-order accurate central finite differences. The mixture energy is also updated to include the elastic contribution, which gives
@@
where $\bm{l}$ is the velocity gradient, and for an isotropic Kelvin–Voigt material $\dot{\bm{T}}^{e}=2G\bm{D}^{d}$ [rodriguez2019] are used to derive a time evolution equation for the elastic stresses. Here, the deviatoric component of the velocity gradient is $\bm{D}^{d}=\bm{D}-\mathrm{tr}(\bm{D})\bm{I}/3$. Appending this evolution equation to the five-equation model gives the five-equation model with hypoelasticity
@@
$\frac{\partial\rho\bm{u}}{\partial t}+\nabla\cdot(\rho\bm{u}\otimes\bm{u}+p\bm{I}-\bm{T}^{e})$ $=\nabla\cdot\bm{T}^{v},$
@@
Details of the hypoelastic model implementation can be found in spratt2024.
Fluid–structure interaction with hyperelastic materials is modeled with an evolution equation for the reference map using the reference map technique (RMT) of [kamrin2012]. The gradient of the reference map $\bm{\xi}$, the inverse of the deformation gradient $\bm{F}$,
@@
and the elastic deviatoric contribution to the Cauchy stress tensor $\bm{T}^{e}$ is computed. For a compressible neo-Hookean material model, the elastic deviatoric part of the Cauchy stress tensor is
@@
where $J$ is the determinant of $\bm{b}$ and $\bm{I}$ is the identity tensor and the hyperelastic energy is $e^{e}=G\left(I_{b}-3\right)/2$, where $I_{b}$ is the first invariant of $\bm{b}$. Further details of the hyperelastic model implementation can be found in ctr24_mrj-mcb.
@@ 4.1.7 Chemical reactions and combustion
High-fidelity simulations of reacting flows are critical to designing efficient propulsion systems. Such simulations require models that account for the effects of chemical reactions on the flow. There are two major differences with respect to simulations of inert flow. First, the mixture composition varies locally due to chemical reactions, and as a result, fluid properties, such as viscosity, also change. This effect must be accounted for to achieve acceptable accuracy; it is typically accomplished using transport equations for individual species. Second, because chemical reactions release heat, temperature changes can become substantial; therefore, variable thermodynamic properties are necessary for accuracy. As we expand on this, we incorporate these effects through detailed models in the gas phase.
@@
Next, we provide detailed expressions for the chemical source term. Their implementation uses code generation and is discussed in section 4.3.7. We simulate gas-phase combustion by integrating section 4.1.7 along with the equations for conservation of momentum and total energy density section 2.1.1. Plans to extend this capability to multi-phase combustion with liquid or solid fuels entail the implementation of detailed expressions for phase changes [ref:Aier2020, ref:Ren2021, ref:Zeng2002].
@@
where $A_{n},\,b_{n},\,T_{a,n}$ are the pre-exponential factor, temperature exponent, and activation temperature of the $n^{\mathrm{th}}$ reaction. The equilibrium constant in 15 is determined from standard equilibrium thermodynamics; detailed expressions can be found elsewhere [ref:cisneros2022, ref:alkhateeb2009].
As stated, one must consider the effect of the mixture thermodynamics on combustion. For the equation of state, we assume a mixture of ideal gases, so
@@
where $\{\hat{h}_{m}(T)\}_{m=1}^{N_{\mathrm{species}}}$ are the species molar enthalpies, modeled using NASA polynomials [ref:mcbride2002]:
@@
with fit coefficients $\{\hat{C}\}_{r=0}^{5}$ determined from experiments and collision integrals [ref:Gardiner1984]. The fitting procedure for 21 includes heats of formation, so, in concert with 19 and 20, it accounts for combustion heat release. In practice, to obtain the temperature, 19 is solved iteratively using a standard Newton method.
The order accuracy is verified via a test problem: the advection of a two-species, calorically perfect mixture. The flow is initialized with a sinusoidal density and temperature profile, as well as spatially uniform species mass fractions. The errors compared to a reference solution are in fig. 7, which shows the expected order of accuracy for all numerical methods.
Figure 7: Convergence results for a 1D 2-component advection problem with constant specific heats.
To validate our implementation, we simulate a one-dimensional reacting shock tube and reproduce the published results of martinez2014. We uniformly discretize the domain of size $L=$12\text{\,}\mathrm{cm}\text{/}$\$ with 400 grid cells. The left boundary represents a reflecting wall, while the right boundary is an outflow. Reactions are modeled using the San Diego mechanism [ref:Saxena2006]. The initial shock is left-going, reflects from the wall, and ignites the mixture. The combustion wave and the shock coalesce into a detonation wave. The comparison between our implementation and the results of martinez2014 is shown in fig. 8. The agreement is sufficient for validation, considering the uncertainties associated with different numerical methods and combustion mechanisms.
Figure 8: (a) Temperature and (b) velocity profiles of a one-dimensional reactive shock tube compared to the results of martinez2014.
@@ 4.1.8 Surface tension at diffuse material interfaces
Surface tension for two-fluid flows is implemented using the model of schmidmayer2017. An advection equation for the color function $c$, which is 0 in one fluid and $1$ in the other, is added to the system of equations. This color function is then used to calculate the capillary stress tensor
@@
$\frac{\partial\rho\bm{u}}{\partial t}+\nabla\cdot(\rho\bm{u}\otimes\bm{u}+p\bm{I}+\bm{\Omega}-\bm{T}^{v})$ $=0,$
@@
where $\varepsilon_{0}$ is the capillary mixture energy. The six-equation model with surface tension is conservative, except for the pressure relaxation terms, and follows the second law of thermodynamics. The implementation of surface tension is validated by resolving the Laplace pressure jump inside a liquid water column in air. The theoretical pressure jump in two dimensions is given by $\Delta P=\sigma/R_{\text{d}}$ where $R_{\text{d}}$ is the radius of the water column. Figure 9 shows the normalized pressure residual
$\varepsilon=\max\left[\frac{\lvert P_{i,j}^{N}-P_{i,j}^{N-1}\rvert}{P_{i,j}^{N}}\right]$
as a function of time and the pressure across the centerline for an $R^{*}=1$ water column with a surface tension coefficient $\sigma^{*}=7.2\times 10^{-4}$ for a $128\times 128$ grid. The pressure jump is resolved exactly, and the residual is time-stable.
Figure 9: Validation results for the surface tension implementation using a Laplace pressure jump problem with a $128\times 128$ grid. (a) shows the normalized residual as a function of time; (b) shows the pressure along the centerline for an $R^{*}=1$ water column with a surface tension coefficient $\sigma^{*}=7.2\times 10^{-4}$.
@@ 4.1.9 Magnetohydrodynamics
Magnetohydrodynamics (MHD) governs the behavior of conducting fluids under the influence of magnetic fields. The governing equations for ideal MHD are [dai1994extension]:
@@
Here, $p$ is the thermal pressure and $\mathbf{B}$ is the magnetic field. The total energy density is $E=\rho e+(\rho|\mathbf{u}|^{2}+|\mathbf{B}|^{2})/2$.
@@
A one-dimensional Brio–Wu MHD shock-tube validation of this implementation is shown in fig. 10 (a), which closely matches the reference [miyoshi2005multi]. The convergence test in fig. 11 also shows the expected order of accuracy for the smooth Alfvén-wave problem [toth2000], confirming the robustness of the implementation.
@@ 4.1.10 Relativistic Magnetohydrodynamics
$\mathbf{U}=\begin{pmatrix}D\\[2.5pt] \mathbf{m}\\[2.5pt] \tau\\[2.5pt] \mathbf{B}\end{pmatrix}=\begin{pmatrix}\Gamma\rho\\[2.5pt] \Gamma^{2}\rho h\mathbf{u}+|\mathbf{B}|^{2}\mathbf{u}-(\mathbf{u}\cdot\mathbf{B})\mathbf{B}\\[2.5pt] \Gamma^{2}\rho h-p+\left[|\mathbf{B}|^{2}+|\mathbf{u}|^{2}|\mathbf{B}|^{2}-(\mathbf{B}\cdot\mathbf{u})^{2}\right]/2-\Gamma\rho\\[2.5pt] \mathbf{B}\end{pmatrix}$.
Here, $\rho$ denotes the rest-mass density, $\mathbf{u}$ is the spatial 3-velocity, $\Gamma$ is the Lorentz factor, $p$ is the thermal pressure, and $h=1+e+p/\rho$ is the specific enthalpy. We adopt natural units, where the speed of light is $c$ and defined by reference as $c=1$. The rest-mass contribution is subtracted from the total energy density to reduce numerical errors in the non-relativistic limit.
The governing equations are [mignone2006hllc]:
@@
$\frac{\partial\mathbf{m}}{\partial t}+\nabla\cdot\left[\mathbf{m}\otimes\mathbf{u}+\left(p+\frac{1}{2}\left(\frac{|\mathbf{B}|^{2}}{\Gamma^{2}}+(\mathbf{u}\cdot\mathbf{B})^{2}\right)\right)\mathbf{I}-\frac{\mathbf{B}\otimes\mathbf{B}}{\Gamma^{2}}-(\mathbf{u}\cdot\mathbf{B})\mathbf{u}\otimes\mathbf{B}\right]$ $=0,$
@@
Recovery of the primitive variables from these highly nonlinear conservative variables requires an iterative scheme. In this work, we solve a single nonlinear equation using the Newton–Raphson method [mignone2006hllc].
Figure 10: Density profiles for (a) ideal MHD and (b) RMHD shock-tube problems. Panel (a) follows miyoshi2005multi (their Fig. 5), and both solvers align closely with the reference solution. Panel (b) follows mignone2006hllc (their Fig. 3), and shows similarly tight agreement; the RMHD case here uses WENO3 rather than the reference’s second-order scheme. Only density is shown; other variables behave comparably.
Figure 11: Convergence test for the smooth Alfvén‐wave problem [toth2000] using WENO3 and WENO5, showing errors decreasing at the expected rate.
To validate the implementation, we consider a relativistic one-dimensional RMHD shock-tube problem based on mignone2006hllc; the resulting density profile in Figure 10 (b) aligns closely with the reference, as do the other variables (not shown).
@@ 4.2.1 General characteristic boundary conditions
The priority for time-dependent boundary conditions for hyperbolic equations is to ensure a smooth egress of the outgoing artifacts with minimal impact on the interior solution. For this purpose, the non-reflecting characteristic boundary conditions [thompson1990] are used, which perform a characteristic decomposition of the governing equation at the boundary. Ignoring the contribution of the viscous and transverse terms, this is
@@
where $q_{c}$ is the set of conservative variables, $\Lambda_{x}=\operatorname{diag}(\lambda_{j})=\operatorname{diag}(u-c,u,u,u,u+c)$ are the eigenvalues of the Jacobian, and $R_{x}$ and $L_{x}$ are the right and left eigenvectors. The contribution of the characteristics can be further decomposed into incoming and outgoing waves, depending on the sign of the eigenvalue, as
@@
[…]
where $\Delta$ is the grid spacing between the boundary ($q_{p}$) and the ghost point ($q_{p}^{\mathrm{(g)}}$). In practice, $\Delta$ is set to be the grid spacing between the boundary and the first interior point for accuracy and stability. For a subsonic outflow at the right boundary, coefficients $\mathcal{L}^{i}_{2}$ through $\mathcal{L}^{i}_{5}$ are set to 0 as these correspond to outgoing waves. The incoming characteristic $\mathcal{L}^{i}_{1}$ is calculated as
@@
One can extend the 1D approach implemented here to treat strong oblique waves. Roughly, one would add a dependence on tangential wavenumbers at the boundary, following the oblique non-reflecting formulations of Giles1990 and the 3D NSCBC extensions by LodatoDomingoVervisch2008. Impedance in the time domain can be enforced via a stable filter [TamDong1996] and time-domain impedance [Jaensch2016], resulting in a non-local time closure that is efficient.